OpenStudy (anonymous):

Need to find bounds for a double polar integral. HALP. Find the area of the surface which is the portion of the cone z = 2sqrt(x^2 + y^2) inside the cylinder x^2 + y^2 = 4.

OpenStudy (anonymous):

Is this correct? \[\int\limits_{0}^{2\pi} \int\limits_{0}^{2} 2r^2 dr d \theta \]